how do i find the new amount of this number increase 14 by 45%
Found 2 solutions by anne, checkley71:
Answer by anne(47)

HI!
i have a math software and it tells that the answer is 20.3 so the increse is 6.3
CHEERS
'anne'

Answer by checkley71(8403)

14+(.45*14)=14+6.3=20.3 ANSWER.